solve the equation

-0.02y + 0.08(10,000-y)=0.15y

Answer by checkley71(8403)

-0.02Y+0.08(10,000-Y)=0.15Y
-0.02Y+800-0.08Y=0.15Y
-0.02Y-0.08Y-0.15Y=-800
-0.25Y=-800
Y=-800/-0.25
Y=3,200 ANSWER.